# Parallelizing the Approximate Minimum Degree Ordering Algorithm: Strategies and Evaluation

### v1

@@ 2.5.2. Approximate Minimum Degree and its Variants
Since the degree updates have been identified as the primary bottleneck of the minimum degree algorithm in the literature (Liu, 1985; George and Liu, 1989), several researchers (Davis and Duff, 1997, 1991; Gilbert et al., 1992; Amestoy et al., 1996) proposed computing upper bounds on the degrees rather than computing exact values. Among these approaches, the AMD algorithm (Amestoy et al., 1996) has been shown to be the most effective. Its sequential implementation in SuiteSparse (Davis, 2019) is now widely used in modern sparse solvers. The success of the AMD algorithm also spurred further interest in nonsymmetric matrices, leading to the development of the column approximate minimum degree algorithm (Davis et al., 2004), which aims to reorder columns to minimize worst-case fill-in across all potential numerical row pivoting during factorization.
2.5.3. Theoretical Development.
@@
The NP-completeness of the fill-in minimization problem was established by Yannakakis (Yannakakis, 1981). Heggernes et al. (Heggernes et al., 2002) analyzed the asymptotic complexity of the minimum degree and AMD algorithms. Additionally, several theoretical developments have been made to understand the minimum degree algorithm and its variants more, including improvements to the minimum degree algorithm, randomized approaches, approximation approaches, and parameterized complexity (Cummings et al., 2021; Fahrbach et al., 2018; Natanzon et al., 1998; Bliznets et al., 2020; Cao and Sandeep, 2020; Fomin and Villanger, 2013).
2.5.4. Tie-breakings.
@@
The importance of tie-breaking strategies in the minimum degree algorithm and its variants has long been recognized by researchers (George and Liu, 1989; Duff et al., 2017, 1976; Amestoy et al., 1996; Cavers, 1987). The quality of the resulting ordering is remarkably sensitive to the initial ordering of the input: changes of the initial ordering can increase or decrease the number of fill-ins by up to a factor of three empirically (George and Liu, 1989). Despite this sensitivity, a consistent effective tie-breaking strategy remains elusive. The state-of-the-art AMD implementation in SuiteSparse (Davis, 2019) still suffers from tie-breaking limitations. To decouple this issue when evaluating reordering algorithms, prior works (George and Liu, 1989; Amestoy et al., 1996) resorted to randomly permuting the input matrix before applying reorderings.
2.5.5. Parallelization.
@@
Unlike nested dissection (George, 1973; Gilbert and Tarjan, 1986), another widely used fill-reducing algorithm, where extensive efforts have been made to develop parallel implementations for both shared and distributed memory systems (Chevalier and Pellegrini, 2008; Karypis and Kumar, 1998; LaSalle and Karypis, 2015), the AMD algorithm has seen little progress in this regard. To our knowledge, only two parallelization attempts related to the AMD algorithm exist. The first is in PT-Scotch (Chevalier and Pellegrini, 2008), where parallel nested dissection is followed by a variant of the AMD algorithm (Pellegrini et al., 2000) independently within each partition; however, the AMD step itself remains sequential and this paradigm relies on the existence of good separators. The second attempt is a parallelization of the column approximate minimum degree algorithm (Chen et al., 1999), in which the minimum degree criterion was removed to increase parallelism across an independent set of pivots. Nevertheless, it resulted in poor scaling and did not report the quality of the resulting orderings. Overall, to the best of our knowledge, there is currently no scalable parallel implementation of the AMD algorithm.
3. Parallel Approximate Minimum Degree Algorithm
@@
In this section, we first explain why simply parallelizing the most computationally expensive component, approximate degree updates in Algorithm 1, is insufficient for achieveing a scalable implementation. We then describe how we address this challenge by adapting the multiple elimination strategy (Liu, 1985) to suit the approximate minimum degree (AMD) algorithm (Amestoy et al., 1996) using distance-2 independent sets. By introducing this new parallel framework and minimizing contention among threads via specialized concurrent data structures tailored to the AMD algorithm, we present the first scalable shared memory AMD implementation.
3.1. Intra-Elimination Parallelism
@@
A straightforward approach to extract parallelism for approximate degree updates from Algorithm 1 is to parallelize by variables in line 1 and use atomics for each access to $w(e)$ between line 1 and line 3 within each elimination step. However, this strategy does not scale. Table 2 shows the average values of $|\mathcal{L}_{p}|$, $\sum_{v\in\mathcal{L}_{p}}|\mathcal{E}_{v}|$, and $\left|\bigcup_{v\in\mathcal{L}_{p}}\mathcal{E}_{v}\right|$ across all elimination steps. While the amount of parallelism, $|\mathcal{L}_{p}|$, appears reasonable, this method is not scalable since the amount of work, $\sum_{v\in\mathcal{L}_{p}}|\mathcal{E}_{v}|$, is not large enough compared to $|\mathcal{L}_{p}|$. This is also confirmed by the observation that each elimination step only takes around tens of microseconds on average. Moreover, contention is high during memory access because the number of unique elements accessed, $\left|\bigcup_{v\in\mathcal{L}_{p}}\mathcal{E}_{v}\right|$, is significantly smaller than the amount of parallelism, $|\mathcal{L}_{p}|$. These factors together illustrate why simply parallelizing the approximate degree updates fails to deliver a scalable implementation.
@@
Table 2. Average sizes of the sets across all elimination steps, representing the amount of parallelism, the amount of work, and the number of unique elements accessed, respectively, by applying intra-elimination parallelism on Algorithm 1.
Matrix Name $|\mathcal{L}_{p}|$ $\sum_{v\in\mathcal{L}_{p}}|\mathcal{E}_{v}|$ $\left|\bigcup_{v\in\mathcal{L}_{p}}\mathcal{E}_{v}\right|$
@@
3.2. Inter-Elimination Parallelism via Distance-2 Independent Sets
@@
Since extracting parallelism within a single elimination step has proven unscalable, we now turn to the possibility of parallelizing the AMD algorithm across elimination steps using multiple elimination (Liu, 1985) discussed in Section 2.3. While the multiple minimum degree algorithm (Liu, 1985) improved sequential performance through multiple elimination, the AMD algorithm did not incorporate this technique. This is primarily because the main advantage of multiple elimination, reducing the cost of degree updates, diminishes when degrees are already approximated (Amestoy et al., 1996). However, multiple elimination provides an additional benefit: it enables parallelism across elimination steps by breaking dependencies between pivot eliminations.
@@
To effectively leverage multiple elimination for parallelizing the AMD algorithm, we need to adapt the policy for selecting the set of pivots. In the original design, performance gains from multiple elimination stemmed from the overlap of pivots’ neighborhoods as explained in Section 2.3. However, such overlap is not well-suited to the AMD algorithm, as the degree of a variable is intended to be approximated with respect to the single adjacent pivot. Moreover, overlapping neighborhoods introduce significant contention during connection updates, since a variable’s updates can be affected by multiple pivots, even when those pivots do not directly interfere with each other. To address these issues, we instead select a distance-2 independent set of pivots, ensuring no overlap in the pivots’ neighborhoods as shown in Figure 4. Pivots in the distance-2 independent sets are then distributed among threads and processed independently in parallel.
@@
Figure 4. When multiple elimination is applied to an independent set $\{1,3,7,9\}$, memory contention arises because multiple pivots attempt to update the connections of shared neighbors—variables 2, 4, 6, and 8. Moreover, since these variables are adjacent to more than one pivot, computing their approximate degrees becomes cumbersome. In contrast, using multiple elimination on a distance-2 independent set $\{1,9\}$ avoids these issues entirely, where variables 2 and 4 are only adjacent to pivot 1 and variables 6 and 8 are only adjacent to pivot 9.
[…]
An example to demonstrate the difference when multiple elimination is applied to a distance-2 independent set instead of an independent set.
@@
We also relax the minimum approximate degree criterion for pivot selection by a multiplicative factor $mult$. Specifically, we select a distance-2 independent set from the candidate variables whose approximate degrees are no greater than $mult$ times the current minimum approximate degree. Table 3 shows the average sizes of maximal distance-2 independent sets across elimination steps for various values of $mult$, where all variables in the set are eliminated in each step. The results show that relaxation is necessary to yield enough pivots to fully utilize all threads when pivots are processed in parallel. However, overly aggressive relaxation can significantly degrade the ordering quality. We will further analyze the trade-off between relaxation factor, ordering quality, and scalability in Section 4.5.
@@
Table 3. Average sizes of maximal distance-2 independent sets across all elimination steps with various values of multiplicative relaxation factor, $mult$, where all variables in the set are eliminated in each step.
Matrix Name $mult=1.0$ $mult=1.1$ $mult=1.2$
@@
3.3. Parallelizing the Core of the Approximate Minimum Degree Algorithm
@@
Once multiple elimination via distance-2 independent sets is applied, most components of the AMD algorithm, including approximate degree updates and the other techniques mentioned in Section 2.4, can be embarrassingly parallelized in shared memory where each thread eliminates a pivot at a time until the distance-2 independent set is depleted. However, two components still suffer from significant contention and dynamic memory management: connection updates and the maintenance of approximate degree lists. In the following, we explain how we overcome these challenges by leveraging specific properties of the AMD algorithm.
3.3.1. Concurrent Connection Updates
@@
In the sequential AMD implementation from SuiteSparse (Davis, 2019), the sets $\mathcal{A},\mathcal{E}$ and $\mathcal{L}$ in the quotient graph are stored in a single compressed sparse column (CSC) structure with additional splitters to track the sizes of $\mathcal{A}_{v}$, $\mathcal{E}_{v}$ and $\mathcal{L}_{e}$ for each variable $v$ and element $e$. As these adjacency sets can grow beyond their initial sizes during connection updates, extra space—referred to as elbow room—is pre-allocated at the end of the CSC structure. When an adjacency set expands, it acquires space from the elbow room and migrates its contents, leaving the old space unused. Once the elbow room is exhausted, garbage collection is triggered to compact the CSC structure and reclaim unused memory, emptying elbow room at the end. Thanks to George and Liu’s result (George and Liu, 1981), which showed that the quotient graph requires no more storage than the original graph, this memory reuse mechanism ensures that no additional memory allocation is needed.
@@
While it is possible to parallelize this structure using atomics to coordinate thread access to the elbow room, the approach suffers during garbage collection, as threads must synchronize and redo their ongoing updates after garbage collection is complete. Instead, we observe empirically that the total extra memory required during elimination never exceeds 1.5 times the storage size of $E^{0}$. This is primarily due to two facts: (i) for all variables $v$, the combined size $|\mathcal{A}_{v}|+|\mathcal{E}_{v}|$ never increases during elimination, and extra memory is only needed when creating $\mathcal{L}_{p}$ for a pivot $p$, whose size does not grow thereafter; and (ii) variables can be merged during elimination using techniques mentioned in Section 2.4, which reduces the number of pivots and, consequently, the amount of extra memory required. Therefore, by augmenting the CSC structure by 1.5 times in advance, we can empirically eliminate the need for garbage collection. To further reduce contention, each thread performs only a single atomic operation to claim extra space after collecting all connection updates associated with the assigned pivots from the distance-2 independent set. Notably, the use of a distance-2 independent set ensures that each variable is associated with at most one pivot, thereby minimizing contention.
@@
[…]
$affinity(v)\leftarrow-1$EndFunction
[…]
3:$dlist_{tid}(loc(v)).remove(v)$EndIf
4:$dlist_{tid}(deg).insert(v)$
5:$loc(v)\leftarrow deg$
6:$affinity(v)\leftarrow tid$EndFunction
[…]
8:$dlist_{tid}(deg).remove(v)$
[…]
10:return entries in $dlist_{tid}(deg)$EndFunction
[…]
Algorithm 3 Parallel Distance-2 Independent Set with known approximate minimum degree $amd$, an analog of Luby’s algorithm (Luby, 1985)
[…]
2:     $Candidates\leftarrow[]$Forinteger $d\in[amd,\lfloor mult\times amd\rfloor$]
[…]
4:     $Candidates.resize(lim)$
[…]
6:     $l(v)\leftarrow(rand(),v)$
7:     $\forall u\in\{v\}\cup N_{v},l_{min}(u)\leftarrow(\infty,\infty)$EndFor
[…]
9:     $\forall u\in\{v\}\cup N_{v}$, atomically update $l_{min}(u)$
10:      by the minimum of $l_{min}(u)$ and $l(v)$EndFor

### v2

@@ Abstract
Since the degree updates have been identified as the primary bottleneck of the minimum degree algorithm in the literature [liu1985modification, george1989evolution], several researchers [davis1997unsymmetric, davis1991unsymmetric, gilbert1992sparse, amestoy1996approximate] proposed computing upper bounds on the degrees rather than computing exact values. Among these approaches, the AMD algorithm [amestoy1996approximate] has been shown to be the most effective. Its sequential implementation in SuiteSparse [suitesparsegithub] is now widely used in modern sparse solvers. The success of the AMD algorithm also spurred further interest in nonsymmetric matrices, leading to the development of the column approximate minimum degree algorithm [davis2004column], which aims to reorder columns to minimize worst-case fill-in across all potential numerical row pivoting.
2.5.3 Theoretical Development.
@@
The fill-in minimization problem was shown to be NP-complete by Yannakakis [yannakakis1981computing]. Heggernes et al. [heggernes2002computational] analyzed the asymptotic complexity of the minimum degree and AMD algorithms. Additionally, several theoretical developments have been made to understand the minimum degree algorithm and its variants more, including improvements to the minimum degree algorithm, randomized approaches, approximation approaches, and parameterized complexity [cummings2021fast, fahrbach2018graph, natanzon1998polynomial, bliznets2020lower, cao2020minimum, fomin2013subexponential].
2.5.4 Tie-breakings.
@@
The need of tie-breaking strategies in the minimum degree algorithm and its variants has long been recognized by researchers [george1989evolution, duff2017direct, duff1976george, amestoy1996approximate, cavers1987tiebreaking]. The ordering quality is remarkably sensitive to the initial ordering of the input: changes of the initial ordering can increase or decrease the number of fill-ins by up to a factor of three empirically [george1989evolution]. Despite this sensitivity, an effective tie-breaking strategy remains elusive. The state-of-the-art AMD implementation in SuiteSparse [suitesparsegithub] still suffers from tie-breaking limitations. To decouple this issue when evaluating reordering algorithms, prior works [george1989evolution, amestoy1996approximate] resorted to randomly permuting the input matrix before applying reorderings.
2.5.5 Parallelization.
@@
In contrast to nested dissection [george1973nested, gilbert1986analysis] where extensive efforts have been made to develop parallel implementations [chevalier2008pt, karypis1998parallel, lasalle2015efficient], the AMD algorithm has seen little progress in this regard. To our knowledge, only two parallelization attempts related to the AMD algorithm exist. The first is in PT-Scotch [chevalier2008pt], where parallel nested dissection is followed by a variant of the AMD algorithm [pellegrini2000hybridizing] independently within each partition; however, the AMD step itself remains sequential and this paradigm relies on the existence of good separators. The second attempt is a parallelization of the column approximate minimum degree algorithm [chen1999toward], in which the minimum degree criterion was removed to increase parallelism across an independent set of pivots. Nevertheless, it resulted in poor scaling and did not report the quality of the resulting orderings. Overall, to the best of our knowledge, there is currently no scalable parallel implementation of the AMD algorithm.
2.5.6 Impact.
While linear systems with multiple right-hand sides or identical sparsity patterns, but varying numerical values, may not gain significant benefits from our approach since the AMD ordering time can be amortized, there are applications where reordering cannot be reused. Specifically, this applies to systems with sparsity patterns that change over time. Examples include Incremental Potential Contact [li2020incremental] and adaptive remeshing [schmidt2023surface, sellan2020opening]. Moreover, the AMD algorithm has shown valuable not only for direct solvers but also for iterative solvers when generating preconditioners [tunnell2025empirical].
3 Parallel Approximate Minimum Degree Algorithm.
@@
In this section, we first explain why simply parallelizing the most computationally expensive component, approximate degree updates in Algorithm 1, is insufficient for achieving a scalable implementation. We then describe how we address this challenge by adapting the multiple elimination strategy [liu1985modification] to suit the approximate minimum degree (AMD) algorithm [amestoy1996approximate] using distance-2 independent sets. By introducing this new parallel framework and minimizing contention among threads via specialized concurrent data structures tailored to the AMD algorithm, we present the first scalable shared memory AMD implementation.
3.1 Intra-Elimination Parallelism.
@@
A simple approach to extract parallelism from Algorithm 1 is to parallelize by variables in line 2 and use atomics for each access to $w(e)$ between line 4 and line 6 within each elimination step. However, this strategy does not scale. As shown in Table 3.1, while the amount of parallelism, $|\mathcal{L}_{p}|$, appears reasonable, this method is not scalable since the amount of work, $\sum_{v\in\mathcal{L}_{p}}|\mathcal{E}_{v}|$, is not large enough compared to $|\mathcal{L}_{p}|$. Moreover, memory contention is high because the number of unique elements accessed, $\left|\bigcup_{v\in\mathcal{L}_{p}}\mathcal{E}_{v}\right|$, is significantly smaller than the amount of parallelism, $|\mathcal{L}_{p}|$. These together illustrate why simply parallelizing the approximate degree updates fails to deliver a scalable implementation.
@@
Table 3.1: Average sizes of the sets across all elimination steps, representing the amount of parallelism $|\mathcal{L}_{p}|$, the amount of work $\sum_{v\in\mathcal{L}_{p}}|\mathcal{E}_{v}|$, and the number of unique elements accessed $\left|\bigcup_{v\in\mathcal{L}_{p}}\mathcal{E}_{v}\right|$, by applying intra-elimination parallelism.
Matrix $|\mathcal{L}_{p}|$ $\sum_{v\in\mathcal{L}_{p}}|\mathcal{E}_{v}|$ $\left|\bigcup_{v\in\mathcal{L}_{p}}\mathcal{E}_{v}\right|$
@@
3.2 Inter-Elimination Parallelism Exploiting Distance-2 Independent Sets.
@@
Since extracting parallelism within a single elimination step has proven unscalable, we now turn to the possibility of parallelizing the AMD algorithm across elimination steps using multiple elimination [liu1985modification] discussed in Section 2.3. While the multiple minimum degree algorithm [liu1985modification] improved sequential performance through multiple elimination, the AMD algorithm did not incorporate this technique [amestoy1996approximate]. This is because the advantage of multiple elimination diminishes when degrees are approximated [amestoy1996approximate]. However, multiple elimination provides an additional benefit: it enables parallelism across elimination steps by breaking dependencies between pivots.
@@
To effectively leverage multiple elimination for parallelizing the AMD algorithm, we need to adapt the policy for selecting pivots. In the original design, performance gains from multiple elimination stemmed from the overlap of pivots’ neighborhoods. However, such overlap is not well-suited to the AMD algorithm, as the degree of a variable is intended to be approximated with respect to the single adjacent pivot. Moreover, overlapping neighborhoods introduce significant contention during connection updates, since a variable’s updates can be affected by multiple pivots, even when those pivots do not directly interfere with each other. To address these issues, we instead select a distance-2 independent set of pivots, ensuring no overlap in the pivots’ neighborhoods as shown in Figure 3.1. Pivots in the distance-2 independent sets are then distributed among threads and processed independently in parallel.
@@
Figure 3.1: When multiple elimination is applied to an independent set $\{1,3,7,9\}$, memory contention arises because multiple pivots attempt to update the connections of variables 2, 4, 6, and 8. In contrast, using multiple elimination on a distance-2 independent set $\{1,9\}$ avoids these issues entirely, where variables 2 and 4 are only adjacent to pivot 1 and variables 6 and 8 are only adjacent to pivot 9.
@@
We also relax the minimum approximate degree criterion for pivot selection by a multiplicative factor $\mathit{mult}$. Specifically, we select a distance-2 independent set from the candidate variables whose approximate degrees are no greater than $\mathit{mult}$ times the current minimum approximate degree. Table 3.2 shows the average sizes of maximal distance-2 independent sets across elimination steps for various values of $\mathit{mult}$, where all variables in the set are eliminated in each step. The results show that relaxation is necessary to yield enough pivots to fully utilize all threads when pivots are processed in parallel. However, overly aggressive relaxation can significantly degrade the ordering quality. We will further analyze the trade-off between relaxation factor, ordering quality, and scalability in Section 4.5.
@@
Table 3.2: Average sizes of maximal distance-2 independent sets with various values of relaxation factor, $\mathit{mult}$.
Matrix $\mathit{mult}=1.0$ $\mathit{mult}=1.1$ $\mathit{mult}=1.2$
@@
3.3 Parallelizing the Core of the Approximate Minimum Degree Algorithm.
@@
Once multiple elimination via distance-2 independent sets is applied, most components of the AMD algorithm, including approximate degree updates and the other techniques mentioned in Section 2.4, can be embarrassingly parallelized where each thread eliminates a pivot at a time until the distance-2 independent set is depleted. However, two components still suffer from significant contention and dynamic memory management: connection updates and the maintenance of approximate degree lists. In the following, we explain how we overcome these challenges.
3.3.1 Concurrent Connection Updates.
@@
In the AMD implementation from SuiteSparse [suitesparsegithub], the sets $\mathcal{A},\mathcal{E}$ and $\mathcal{L}$ in the quotient graph are stored in a single compressed sparse column (CSC) structure with additional splitters to track the sizes of $\mathcal{A}_{v}$, $\mathcal{E}_{v}$ and $\mathcal{L}_{e}$ for each variable $v$ and element $e$. As these adjacency sets can grow beyond their initial sizes during connection updates, extra space referred to as elbow room is pre-allocated at the end of the CSC structure. When an adjacency set expands, it acquires space from the elbow room and migrates its contents, leaving the old space unused. Once the elbow room is exhausted, garbage collection is triggered to compact the CSC structure and reclaim unused memory, emptying elbow room at the end.
@@
While it is possible to parallelize this structure using atomics to coordinate thread access to the elbow room, the approach suffers during garbage collection, as threads must synchronize and redo their ongoing updates after garbage collection is complete. Instead, we observe empirically that the total extra memory required during elimination never exceeds 1.5 times the storage size of $E^{0}$. This is primarily due to two facts: (i) for all variables $v$, the combined size $|\mathcal{A}_{v}|+|\mathcal{E}_{v}|$ never increases during elimination, and extra memory is only needed when creating $\mathcal{L}_{p}$ for a pivot $p$, whose size does not grow thereafter; and (ii) variables can be merged during elimination using techniques mentioned in Section 2.4, which reduces the number of pivots and, consequently, the amount of extra memory required. Therefore, by augmenting the CSC structure by 1.5 times in advance, we can empirically eliminate the need for garbage collection. We acknowledge that some inputs may not follow this empirical observation, so we allow users to adjust the augmentation factor when necessary. To further reduce contention, each thread performs only a single atomic operation to claim extra space after collecting all connection updates associated with the assigned pivots from the distance-2 independent set. Notably, the use of a distance-2 independent set ensures that each variable is associated with at most one pivot, thereby minimizing contention.
3.3.2 Concurrent Lists for Maintaining Approximate Degrees.
In the AMD implementation from SuiteSparse [suitesparsegithub], pivot selection is facilitated by maintaining $n$ degree lists. Each list corresponds to a specific approximate degree and stores the associated variables using a doubly linked list. When a pivot is selected, it is removed from the list, and for each variable in its neighborhood, it is moved from its current list to the new one based on its updated approximate degree. No dynamic memory allocation is required, as the total number of entries across all lists is bounded by $n$. However, this data structure is not amenable to parallelization due to high contention when multiple threads attempt to access the same doubly linked list.
Instead, we let each thread $\mathit{tid}$ maintain its own structures: $n$ doubly linked lists $\mathit{dlist_{tid}}(0,\cdots,n-1)$ to store variables based on their approximate degrees, a $\mathit{loc}$ array to record the local degree list to which each variable belongs, and an $\mathit{lamd}$ value to store the minimum approximate degree of variables locally maintained by this thread. In addition, a shared $\mathit{affinity}$ array is used to track which thread holds the most up-to-date information for each variable. Both the $\mathit{affinity}$ and $\mathit{loc}$ arrays are initialized to 1-1, and $\mathit{lamd}$ is initialized to $n$. The use of a distance-2 independent set of pivots ensures that each variable is associated with at most one pivot—and thus with a single thread—during any given elimination step.
@@
1: function Remove($tid,v$)
2:   $\mathit{affinity}(v)\leftarrow-1$
3:
4: function Insert($tid,v,deg$)
5:   if $loc(v)\neq-1$ then
6:    $dlist_{tid}(loc(v)).remove(v)$
7:   $dlist_{tid}(deg).insert(v)$
8:   $loc(v)\leftarrow deg$
9:   $\mathit{affinity}(v)\leftarrow tid$
10:   $lamd\leftarrow\min\{deg,lamd\}$
11:
12: function Get($tid,deg$)
13:   for $v\in dlist_{tid}(deg)$ do
14:    if $\mathit{affinity}(v)\neq tid$ then
15:      $dlist_{tid}(deg).remove(v)$
16:      $loc(v)\leftarrow-1$
17:   return entries in $dlist_{tid}(deg)$
18:
19: function Lamd(tidtid)
20:   while $lamd<n$ and $\textsc{Get}(tid,lamd)=\emptyset$ do
21:    $lamd\leftarrow lamd+1$
22:   return lamdlamd
As shown in Algorithm 1, when a thread attempts to remove a variable $v$, it sets $\mathit{affinity}(v)$ to 1-1, effectively invalidating all entries of $v$ across all threads without modifying the lists. When a thread attempts to insert a variable $v$, it first checks if $v$ is currently stored in the local degree lists based on $\mathit{loc}(v)$. If so, i.e. $\mathit{loc}(v)\neq-1$, the stale entry of $v$ in the old degree list is explicitly removed. In all cases, the variable $v$ is then inserted into the new degree list in thread $\mathit{tid}$, with $\mathit{loc}(v)$ set to the new approximate degree and $\mathit{affinity}(v)$ set to $\mathit{tid}$, indicating that the thread $\mathit{tid}$ has the latest information for variable $v$. The value of $\mathit{lamd}$ is also updated accordingly. Stale entries of variable $v$ stored in other threads, identified by a mismatch between $\mathit{affinity}(v)$ and the thread ID, are lazily reclaimed during list traversal for pivot selection using the Get function. Since the value of $\mathit{lamd}$ is not updated when removing variables, an Lamd function is used to skip the empty degree lists and get the up-to-date minimum approximate degree maintained by a thread when requested. This design avoids contention for degree list maintenance, with the only necessary coordination occurring when determining the global minimum approximate degree.
Algorithm 2 Parallel Distance-2 Independent Set, an analog of Luby’s algorithm [luby1985simple]
1: function Dist_2_Indep_Set($mult,lim$)
2:   $amd\leftarrow\min_{tid}(\textsc{Lamd}(tid))$
3:   parallel for each thread $\mathit{tid}$ do
4:    $Candidates\leftarrow[]$
5:    for integer $d\in[amd,\lfloor mult\times amd\rfloor$] do
6:      $Candidates.append(\textsc{Get}(tid,d))$
7:      if $Candidates.size()\geq lim$ then
8:       $Candidates.resize(lim)$
9:       break
10:    for variable $v\in Candidates$ do
11:      $l(v)\leftarrow(rand(),v)$
12:      $\forall u\in\{v\}\cup N_{v},l_{min}(u)\leftarrow(\infty,\infty)$
13:    barrier
14:    for variable $v\in Candidates$ do
15:      $\forall u\in\{v\}\cup N_{v}$, atomically update $l_{min}(u)$
16:      by the minimum of $l_{min}(u)$ and $l(v)$
17:    barrier
18:    for variable $v\in Candidates$ do
19:      if $l(v)=l_{min}(u),\forall u\in\{v\}\cup N_{v}$ then
20:       $v$ is a valid candidate
21:   return all valid candidates
3.4 Parallel Distance-2 Independent Set.
To parallelize the selection of a distance-2 independent set for multiple elimination, we first compute the global minimum approximate degree, $\mathit{amd}$, from the concurrent approximate degree lists. We then collect candidate variables whose approximate degrees fall within the relaxed threshold described in Section 3.2. This collection is performed in parallel from the local degree lists maintained by each thread. To limit overhead compared to the sequential AMD algorithm, we restrict each thread to collect at most $\mathit{lim}$ candidates. The impact of this parameter will be analyzed in Section 4.5.
To select a distance-2 independent set, we perform a single iteration of the distance-2 analog of Luby’s algorithm [luby1985simple] as shown in Algorithm 2. For each candidate gathered from the approximate degree lists, a random number is generated and compared against the candidates within its distance-2 neighborhood, with ties broken by candidate indices. A candidate is considered valid if and only if it has the smallest value among the candidates within its distance-2 neighborhood. To enable parallelism, Algorithm 2 includes appropriate atomic operations and synchronization barriers. We intentionally avoid running the full algorithm because a maximal distance-2 independent set is unnecessary; as long as the set obtained in the first iteration is sufficiently large, it provides enough parallelism for the core AMD algorithm.
3.5 Overall Parallel AMD Algorithm.
Algorithm 3 summarizes our strategies used to parallelize the AMD algorithm. These techniques collectively enable an efficient and scalable implementation, addressing challenges in parallelizing the AMD algorithm.
Algorithm 3 Parallel AMD Algorithm
1: while $|V|>0$ do
2:   $D\leftarrow\textsc{Dist\_2\_Indep\_Set}(mult,lim)$
3:   parallel for pivot $p$ in $D$ do
4:    Eliminate $p$ using the core AMD algorithm
5:    with concurrent connection updates and
6:    concurrent approximate degree lists
7:    (see Section 3.3)
3.5.1 Memory Consumption.
Let $n$ be the number of rows, $m$ the number of nonzeros in the input matrix, and $t$ the number of threads. The AMD implementation from SuiteSparse [suitesparsegithub] requires $O(n+m)$ memory. On top of that, our implementation allocates extra $1.5m$ memory for CSC augmentation, $O(nt)$ memory for the concurrent approximate degree lists, and $O(n)$ memory for distance-2 independent set selection. All big-$O$ terms involve small constants, and in practice $nt$ is usually smaller than, or on the same order as, $m$; therefore, we expect our method to incur only a modest constant factor increase in memory.
4 Experiments.
4.1 Evaluation Platforms.
We evaluated our parallel AMD algorithm [^1] https://github.com/PASSIONLab/ParAMD on an AMD EPYC 7763 CPU with 64 cores. The implementation was written in C++ using OpenMP and compiled with GCC 13.2.1 using the flags -O3 -fopenmp. We set OMP_PROC_BIND=close so threads are in the same NUMA domain when possible. As a baseline, we used the sequential implementation from SuiteSparse (v7.12.1) [suitesparsegithub]. To the best of our knowledge, no existing parallel implementation of the AMD algorithm is available for comparison.
4.2 Matrix Suite.
The sparse matrices used in our experiments were selected from the SuiteSparse Matrix Collection [davis2011university] and the M3E Matrix Collection [m3e], covering a diverse range of large-scale linear systems arising from real-world applications. A summary of these matrices is provided in Table 4.1.
Note that the AMD algorithm remains a viable option for nonsymmetric solvers such as UMFPACK [davis2004algorithm] when the nonsymmetric matrix exhibits a nearly symmetric nonzero pattern. For this reason, we also included several nonsymmetric matrices in our dataset. In such cases, the AMD algorithm is applied to the matrix obtained by summing the absolute values of the original matrix and its transpose. This pre-processing step is performed by SuiteSparse AMD regardless of whether the input matrix is symmetric. For fair comparison, we parallelized this step using simple atomic operations and included its runtime in the reported results. However, users may skip this pre-processing step in our implementation when the input matrix is known to be symmetric.
Table 4.1: A summary of the matrices used in our experiments, sorted by #nonzeros.
Matrix #rows #nonzeros Symmetric Positive-definite Description
nd24k 72.0K 28.7M $\checkmark$ $\checkmark$ 3D mesh problem
ldoor 952K 42.5M $\checkmark$ $\checkmark$ Structural problem
Serena 1.39M 64.5M $\checkmark$ $\checkmark$ Structural problem
dielFilterV3real 1.10M 89.3M $\checkmark$ $\times$ Electromagnetic problem
ML_Geer 1.50M  [^1] M $\times$ $\times$ Poroelastic problem
Flan_1565 1.56M 114M $\checkmark$ $\checkmark$ Structural problem
Cube_Coup_dt0 2.16M 124M $\checkmark$ $\times$ Structural problem
Bump_2911 2.91M 128M $\checkmark$ $\checkmark$ Geomechanical problem
Cube5317k 5.32M 223M $\checkmark$ $\checkmark$ Elasticity problem
HV15R 2.02M 283M $\times$ $\times$ Fluid dynamics problem
Queen_4147 4.15M 317M $\checkmark$ $\checkmark$ Structural problem
stokes 11.4M 349M $\times$ $\times$ Semiconductor process problem
guenda11m 11.5M 512M $\checkmark$ $\checkmark$ Geomechanical problem
agg14m 14.1M 633M $\checkmark$ $\checkmark$ Mesoscale problem
rtanis44m 44.8M 748M $\checkmark$ $\checkmark$ Diffusion problem
nlpkkt240 28.0M 761M $\checkmark$ $\times$ Optimization problem
4.3 Ordering Comparison.
We now present the performance of our parallel AMD implementation compared to the SuiteSparse baseline, as summarized in Table 4.2. We used the default relaxation factor $\mathit{mult}=1.1$ and set the limitation factor $\mathit{lim}$ to $8192$ divided by the number of threads, following the heuristic that this configuration targets a fill-in ratio of $1.1\times$ over the SuiteSparse baseline while ensuring a sufficiently large pool of candidates is considered for the distance-2 independent set in each elimination step.
Table 4.2 shows that our parallel implementation achieves up to a $7.29\times$ speedup over SuiteSparse using 64 threads. Also, our fill-in ratio is consistently around $1.1\times$ over SuiteSparse, reflecting the effect of the relaxation factor $\mathit{mult}$ in avoiding the selection of overly suboptimal pivots.
Table 4.2: Ordering comparison between SuiteSparse AMD and our 64-thread parallel AMD implementation. Matrices were randomly permuted five times with all methods evaluated on the same set of permutations to decouple tie-breaking issue explained in Section 2.5.4. Ordering time is reported as mean $\pm$ standard deviation due to its variability across runs, while the fill-in results are more consistent where the standard deviation is at most 6% of the mean so they are reported as mean only over the five runs. Speedup is computed per run and we report the mean speedup across five runs.
Matrix Ordering Time (sec) Our 64-thread #Fill-ins Our
SuiteSparse Ours Speedup over SuiteSparse Ours Fill-in
SuiteSparse Ratio